HETAV PATEL

Grade 11,

What weight of CaCO3 shall be neutralised by 100 ml of 0.1 M(molarity) HCl?

What weight of CaCO3 shall be neutralised by 100 ml of 0.1 M(molarity) HCl?

Grade:9

2 Answers

Mahima Kanawat
6 years ago
For neutralisation , No. Of equivalent of CaCO3 = no. Of equivalent of HClNo. Of equivalent of HCL = normality × volume =molarity × valency factor × volume = 0.1 × 1 × 0.1 = 0.01 Now equivalent of calcium carbonate should be equal to equivalent of HCL0.01= wt /equivalent wt ( equivalent wt = molecular wt / valency factor ) = 100 /2 = 50 0.01 = wt / 50Wt = 0.5 gm
Ankit
6 years ago
No of equvelence of HCl = N×M×V = 0.01No. of equvelence of CaCO3= (wt taken) /equvelence wtequvelence wt of CaCO3= molecular wt/